The question asks to find the sum
8+14+20+...+278

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

some of an arithmetic series
Sn=(n/2)(2a+(n-1)d)
a=8
d=6
the range is 270, so n=46, n-1=45
=23(16+270)
=23(286)=6578

Answer by ikleyn(52788)

.

For introductory lessons on arithmetic progressions see
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Arithmetic progressions".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.